A particle p1 is projected vertically upwards from horizontal ground, with a speed of 30m/s.At the same constant another particle p2 is projected vertically upwards from the top of a tower of height 25m,with a speed of 10m/s. Find the time for which p1 is higher than the top of the tower
1
Expert's answer
2019-06-04T12:13:11-0400
"h_1=v_1t-0.5gt^2"

"h_1-h=v_1t-0.5gt^2-h=30t-5t^2-25"

"30t-5t^2-25=0""t_1=1\ s"

"t_2=5\ s"

p1 is higher than the top of the tower between these values. The related time is


"\Delta t=t_2-t_1=5-1=4\ s"